# Long-term Causal Inference via Modeling Sequential Latent Confounding

### v1

@@ 7 Conclusion
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
@@
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)|\mathbf{X},G=O,A=1\right]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)|\mathbf{X},G=O,A=1\right]\}\times p(A=1|\mathbf{X},G=O),$
@@
$\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]$ (16)
$=$ $\mathbb{E}[S|\mathbf{X},G=O,A=1]-\mathbb{E}[S|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)|\mathbf{X},G=O,A=1]\}\times p(A=0|\mathbf{X},G=O)$
+$+\{\mathbb{E}[S(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)|\mathbf{X},G=O,A=1]\}\times p(A=1|\mathbf{X},G=O)$
@@
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=O]-\mathbb{E}[S(0)|\mathbf{X},G=O]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)|\mathbf{X},G=E]-\mathbb{E}[S(0)|\mathbf{X},G=E]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S|\mathbf{X},G=E,A=1]-\mathbb{E}[S|\mathbf{X},G=E,A=0]-\mathbb{E}[S|\mathbf{X},G=O,A=1]+\mathbb{E}[S|\mathbf{X},G=O,A=0]$
@@
Appendix B Proof of Proposition 4.3
@@
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (19)
=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O]$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
@@
$\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]$ (20)
$=$ $\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
@@
$\iff$ $\mathbb{E}[S_{t}(1)|\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)|\mathbf{X},G=E]-\mathbb{E}[S_{t}|\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}|\mathbf{X},G=O,A=0]$
@@
Appendix C Proof of Theorem 4.5
@@
$=$ $\mathbb{E}[Y(1)-Y(0)|\mathbf{X}=\mathbf{x}]$
@@
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (23)
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]p(A=1|\mathbf{X},G=O)+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]p(A=0|\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]p(A=1|\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
@@
where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)|\mathbf{X}$, and last equality is based on Assumption 4.1.
@@
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)|\mathbf{X},G=O]-\mathbb{E}[Y(0)|\mathbf{X},G=O]$ (24)
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]$
@@
$\mathbb{E}[S_{T}(1)|\mathbf{X}]-\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (25)
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)|\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)|\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
@@
$\tau(\mathbf{X})=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)|\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)|\mathbf{X}]$ (26)
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y|\mathbf{X},G=O,A=1]-\mathbb{E}[Y|\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=E,A=0]$
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}|\mathbf{X},G=O,A=0]$
@@
Appendix D Proof of Lemma 5.4
@@
We now restate our Lemma 5.4 as follows and formally prove that.
Lemma D.2.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2 and 5.3 hold, then we have
$\hat{f}(\mathbf{x})-f(\mathbf{x})=O_{p}\left((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}+(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}}+(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\right).$ (29)
@@
We apply Proposition 1 of Kennedy et al. (Kennedy, 2023), yielding that
@@
$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})|\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})|\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$. Here, $f$ is a time-series model ($1$-order autoregressive), optimized with $T-1$-length steps and $n$ samples, as shown in Eq. (8). According to Proposition 4.3, $\forall t$, $f(\mathbf{x})=\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}$. We first prove the rate of $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}$ for a fixed $t$ as follows.

### v2

@@ Appendix D Additional Experimental Details and Results
$=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$
@@
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]\}\times p(A=0\mid\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]\}\times p(A=1\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}\left[S(1)\mid\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(1)\mid\mathbf{X},G=O,A=1\right]\}\times p(A=0\mid\mathbf{X},G=O)$
+$+\{\mathbb{E}\left[S(0)\mid\mathbf{X},G=O,A=0\right]-\mathbb{E}\left[S(0)\mid\mathbf{X},G=O,A=1\right]\}\times p(A=1\mid\mathbf{X},G=O),$
@@
$\mathbb{E}[S(1)\mid\mathbf{X},G=O]-\mathbb{E}[S(0)\mid\mathbf{X},G=O]$ (16)
$=$ $\mathbb{E}[S\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S(1)\mid\mathbf{X},G=O,A=1]\}\times p(A=0\mid\mathbf{X},G=O)$
+$+\{\mathbb{E}[S(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S(0)\mid\mathbf{X},G=O,A=1]\}\times p(A=1\mid\mathbf{X},G=O)$
@@
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)\mid\mathbf{X},G=O]-\mathbb{E}[S(0)\mid\mathbf{X},G=O]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S(1)\mid\mathbf{X},G=E]-\mathbb{E}[S(0)\mid\mathbf{X},G=E]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
$+\mathbb{E}[S\mid\mathbf{X},G=E,A=1]-\mathbb{E}[S\mid\mathbf{X},G=E,A=0]-\mathbb{E}[S\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S\mid\mathbf{X},G=O,A=0]$
@@
Appendix F Proof of Proposition 4.3
@@
$\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]$ (19)
=$\overset{(a)}{=}$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O]$
$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
@@
$\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]$ (20)
$=$ $\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
@@
$\iff$ $\mathbb{E}[S_{t}(1)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}(0)\mid\mathbf{X},G=E]-\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=1]+\mathbb{E}[S_{t}\mid\mathbf{X},G=O,A=0]$
@@
Appendix G Proof of Theorem 4.7
@@
$=$ $\mathbb{E}[Y(1)-Y(0)\mid\mathbf{X}=\mathbf{x}]$
@@
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$ (23)
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]p(A=1\mid\mathbf{X},G=O)+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]p(A=0\mid\mathbf{X},G=O)$
$+\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]p(A=1\mid\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
@@
where the first equality is based on Assumption 3.5, i.e., $G{\perp\!\!\!\perp}Y(a)\mid\mathbf{X}$, and last equality is based on Assumption 4.1.
@@
$\tau(\mathbf{X})=$ $\mathbb{E}[Y(1)\mid\mathbf{X},G=O]-\mathbb{E}[Y(0)\mid\mathbf{X},G=O]$ (24)
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]$
@@
$\mathbb{E}[S_{T}(1)\mid\mathbf{X}]-\mathbb{E}[S_{T}(0)\mid\mathbf{X}]$ (25)
$=$ $\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
+$+\{\mathbb{E}[S_{T}(1)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(1)\mid\mathbf{X},G=O,A=1]\}p(A=0,\mathbf{X},G=O)$
+$+\{\mathbb{E}[S_{T}(0)\mid\mathbf{X},G=O,A=0]-\mathbb{E}[S_{T}(0)\mid\mathbf{X},G=O,A=1]\}p(A=1,\mathbf{X},G=O)$
$=$ $\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
@@
$\tau(\mathbf{X})$ (26)
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(1)\mid\mathbf{X}]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}(0)\mid\mathbf{X}]$
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
$=$ $\mathbb{E}[Y\mid\mathbf{X},G=O,A=1]-\mathbb{E}[Y\mid\mathbf{X},G=O,A=0]+f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=E,A=1]$
$-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=E,A=0]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=1]-f^{\mu}(\mathbf{X})\mathbb{E}[S_{T}\mid\mathbf{X},G=O,A=0]$
@@
Appendix H Proof of Lemma 5.4
@@
We now restate our Lemma 5.4 as follows.
Lemma H.2.
Suppose the training steps S1 and S2 are train on two independent datasets of size $n$ respectively, and suppose Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, 4.1, 5.2, and 5.3 hold, then we have
which attains the oracle rate if $\min\{\alpha,\beta\}\geq\eta$.
@@
We apply Proposition 1 in Kennedy et al. [25], yielding that
@@
$=$ $\hat{\mathbb{E}}_{n}[\hat{r}(\mathbf{X})\mid\mathbf{X}=\mathbf{x}]+o_{p}(R_{n}^{*}(\mathbf{x}))+O_{p}(R_{n}^{*}(\mathbf{x}))$
where $\hat{r}(\mathbf{x})=\mathbb{E}[\hat{f}(\mathbf{X})\mid\mathbf{X}=\mathbf{x}]-f(\mathbf{x})$, and $\mathcal{R}_{n}^{*}(x)$ is the oracle risk of second-stage regression and further under Assumption 5.2, we know $f$ is $\eta$-smooth, thus $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$. Here, $f$ is a time-series model ($1$-order autoregressive), optimized with $T-1$-length steps and $n$ samples, as shown in Eq. (8). According to Proposition 4.3, $\forall t$, $f(\mathbf{x})=\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}$. We first prove the rate of $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}$ for a fixed $t$ as follows.